If one outlet pipe can drain a tank in 24 hours ad another pipe can drain the tank in 26 hours, how long it will take for both pipes to drain the tank?
Answer by Cromlix(4381)

1/24 + 1/26 = 1/x (rate)
Denominator = 624x
26x + 24x = 624
50x = 624
x = 624/50
x = 12.48 hours
x = 12 hours 29mins.
Hope this helps.
:-)